Progress Toward the Milestones of the enclosed schedule:
NOTE: Meetings are held every Tuesday in CEBAF Center,
Room #A110, at 11 a.m., unless otherwise notified.
NOTE: A reminder to all principals to submit a brief weekly progress to D. Gilchrist each Friday by noon
ATTENDEES: G. Biallas, S. Benson, C. Bohn, L. Dillon-Townes,
F. Dylla, D. Engwall, K. Hovater, K. Jordan, G. Krafft,
W. Lacy, R. Legg, G. Neil, C. Rode, T. Seeberger, M. Shinn,
T. Siggins, C. Sinclair, R. Walker, M. Wiseman
During use of the SOP it was realized that the documentation of the Laser PSS was not complete enough. The language has been changed. We will re-certify the PSS at a convenient point.
We will soon be in an operations mode where e adhering to the day to day schedule will be crucial to get our required work done on time. A stand-up meeting of some sort will probably be needed to coordinate the activities.
F. Dylla reminded us that operations time is money and if we should buy the spares, even at the $10000 level.
We should look at our documentation and change labels from 500 kV to the other values as appropriate.
Safety Action Items:
New Safety Action Items:
Resolve the request by the Safety Professionals that the Safety
Officer be appointed for one year rather than one month
Completed Safety Action Items:
T. Siggins/D. Engwall - Get the gun and line baked and get the Experiment running. D. Engwall achieved 300 kV over the week end and quiet running at 250 kV for three hours without a leak. The NEG Pumps were fully activated and a cathode is in the process of being formed. The HeNe laser is set for first beam and the light box optics have been re constituted.
G. Biallas - Make a new ceramic stack and make a new cathode support tube with a anti field emission coating. The schedule was altered to detail the ceramic tube replacement and coating on the cathode support tube activity. We have also added another track of doing Electron Beam Polishing See the weekly report by L. Phillips .
R. Legg - Investigate an alternate scheme to use the Longitudinal apparatus to perform the high voltage portions of the Transverse experiment. B. Dillon Townes will put together a drawing of the proposed line , working from information from D. Kehne. D. Kehne will get the proposal done by Aug. 22. No progress. C. Bohn wants to be at the meeting between D. Kehne & L. Dillon -Townes.
Determine the need for additional personnel required to match or beat the Schedule. No needs this week.
T. Siggins - Write the tricks of polishing learned at SLAC. No action.
G. Biallas - Finish the LCW System: Flow & Temp reading, Level indicators, Full qualification walk through needed. No action.
K. Jordan/S. Benson - Failure analysis for the 10 MeV beam handling. No action
M. Shinn - Put phase noise documentation on the web. The Phase noise system
All - general cleanup after bake. Checkout systems, straighten out cabling. Partially done.
R. Legg - Publish the straw man for the 10 MeV Experiment . Files need to be compatible with the Web.
G. Biallas - integrate the new experimental schedule developed by Bob Legg into the main schedule. In process. No progress this week.
B. Dillon Townes - Get the 10 MeV beam line signed off. The remaining third in signature this week.
K. Jordan - Get the bunch length measurement contract signed. In Procurement's hands
K. Bohn - Produce the design for the multislit for 10 MeV. K. Jordan will be the Engineer. The design is started. K. Jordan will get the schedule bars to G. Biallas.
D. Engwall & D. Kehne - Present the Longitudinal simulation results at 250 kV. D. Engwall will start the final runs tomarrow.
C. Piller - Have both the chopper cavity and the buncher cavity plated or painted to withstand the 250 °C bake. No progress
D. Kehne - In the proposed, combined 500 kV proposal, utilize the a spare bakable solenoid from the injector group for the solenoid in front of the Chopper rather than the built-in solenoid. No progress
K. Bohn/R. Legg - Resolve where the Quartercryomodule should be tested first, in the Cryomodule Test Cave, testing one cavity at a time for basically a window test or to go through the disruption now and test in the ITS Enclosure in the offset position, coincident with the Longitudinal experiment. No progress.
New Action Items:
G. Biallas - Resolve another way to install the Running Resistor. See D. Engwall
T. Siggins - Resolve a way to guarantee the gun or line don't get damaged while anyone works above them and enforce the protection requirement.
T. Siggins - (Added by F. Dylla after the meeting) Do a walk through the gun and beam line and install shields on ceramics and strain relief on cables to prevent the line from being damaged.
S. Benson - Find out the material Duke used to eliminate the corrosion effects of a bake on copper.
L. Dillon-Townes/C. Sinclair/H. Liu - Do the simulations and Resolve if the solenoid before the cryomodule can have its ID made larger by 1 cm.
S. Benson - Get the simulation data from B. Yunn to find out if we need to have a shielded valve between the gun and light box and order the appropriate valve.
Completed Action Items:
Individual Reports of 8/16/96:
WESGO reports that their delivery of the ceramic bodies has slipped to Sept. 20 from Sept. 3. D. Lloyd is working the problem.
The first concept for the rotisserie to do the Ion implantation on the ceramics at LBL was sketched by L. Phillips and will be drawn up over the weekend.
Six coupons of ion implanted alumina were received from LBL and the first was tested with good results.
The inside of the material for the new cathode support tube is being electropolished before machining.
The electrodes to assess the usability of the anti field emission coating for the cathode support tube are being made in the shop for delivery early next week. Tom Venhaus will test them before and after coating in his apparatus at W&M. His apparatus is running.
The new slug of material for the new Mid Voltage Ring is reported
to be to specification at Western Forge & Flange.
Received the rebuilt circulator from Germany, and tested it to full power both in to a matched load and into a short circuit. The circulator performs as required.
Initial checkout of the 50 KW systems is complete. We will
be off for about 2 weeks while we complete items on our punch
lists. we will be ready for final commissioning of the high power
RF in September.
Extensive work occurred this week to identify and resolve problems
with the drive laser's SHG output. A possible interaction between
the rf systems and the drive laser was investigated and found
to not occur. (Thanks to J. Fugitt and M. Augustine's crew.)
I now believe the SHG temperature control loop is failing, and
am awaiting Coherent's suggestions. In the meantime, I can adjust
the temperature if the SHG power output changes.
New drive probes for the Chopper cavity are being designed by C. Piller. It will be several weeks before they will be ready for testing.
Five hundred meters of fiber optic cable for the Arc Detectors on the 50 KW system was received this week and is being installed.
Discussions was held again this week on the RF system software for the IR FEL. A list of parameters was given to the software group.
The wiring is being reworked in the Power Relay chassis for the Buncher cavity water skid controller. It will be ready for testing again early next week.
The front and rear panels for the Arc Detector chassis are due next week. The finished units will be assembled, tested, and tested by the end of the month.
A test of 4 klystrons and 2 circulators operating at 8 KW is
planned for next week.
ITS Schedule for 8/20/96